OpenStudy (anonymous):

simplify radical 5/8?

OpenStudy (anonymous):

\[\sqrt{5/8}\]

OpenStudy (anonymous):



OpenStudy (zpupster):

sqrt5/8= sqrt (5/2*1/4) but 1/2*1/2 = 1/4 so sqrt(5/2)/2 still kind of ugly how abot sqrt 5/8*sqrt2/2=(sqrt 10)/4=